OpenStudy (anonymous):

Solve using the substitution method 8x+3y=-34 -8x+y=42

OpenStudy (anonymous):

x=-5 and y=2

OpenStudy (anonymous):

-8x = 34 + 3y 34 + 3y + y = 42 4y = 8 => y = 2 -8x = 42 -2 => -8x = 40 => x = -5

OpenStudy (anonymous):

-8x+y=42,so y=42+8x because 8x+3y=-34 8x+3(42+8x)=-34 32x=-34-126=-160 x=-5,so y=2

OpenStudy (anonymous):

8x+3y=-34 -8x+y=42 8 cancels and leaves you with 4 y so -34+42 leaves you with 8 so the y answer is definitely 2

OpenStudy (anonymous):

shd101 can you take a look at my question